A recent study of two vendors of desktop personal computers reported that out of 879 units sold by Brand A, 116 required repair, while out of 749 units sold by Brand B, 127 required repair. Round all numeric answers to 4 decimal places. 1. Calculate the difference in the sample proportion for the two brands of computers, ?̂??????−?̂??????= 3. Calculate the pooled estimate of the sample proportion, ?̂= 4. Is the success-failure condition met for this scenario? A. Yes B. No 5. Calculate the test statistic for this hypothesis test. = 6. Calculate the p-value for this hypothesis test, p-value = . 7. Based on the p-value, we have: A. little evidence B. some evidence C. extremely strong evidence D. strong evidence E. very strong evidence that the null model is not a good fit for our observed data. 8. Compute a 90 % confidence interval for the difference ?̂??????−?̂?????? p^BrandA−p^BrandB= ( , )
Answer #1

Answer:

A recent study of two vendors of desktop personal computers reported that out of 879 units sold by Brand A, 116 required repair, while out of 749 units sold by Brand B, 127 required repair. Round all numeric answers to 4 decimal places.

1. Calculate the difference in the sample proportion for the two brands of computers, ?̂??????−?̂??????=

=116/879 – 127/749

=0.1320-0.1696 = -0.0376

3. Calculate the pooled estimate of the sample proportion, ?̂

= (116+127)/(879+749) =0.1493

4. Is the success-failure condition met for this scenario?

A. Yes

5. Calculate the test statistic for this hypothesis test. =

= -2.1214

6. Calculate the p-value for this hypothesis test,

p-value =0.0339 .

7. Based on the p-value, we have:

A. little evidence B. some evidence C. extremely strong evidence D. strong evidence E. very strong evidence that the null model is not a good fit for our observed data.

Note: at 5% and 10% level of significance, there is significant difference between the two brands. Interpret appropriately. At 1% level is not significant. Based on class notes, make your option.

8. Compute a 90 % confidence interval for the difference ?̂??????−?̂?????? p^BrandA−p^BrandB

= (-0.0669, -0.0082 )

Z Test for Differences in Two Proportions

Data

Hypothesized Difference

0

Level of Significance

0.05

Group 1

Number of Items of Interest

116

Sample Size

879

Group 2

Number of Items of Interest

127

Sample Size

749

Intermediate Calculations

Group 1 Proportion

0.1320

Group 2 Proportion

0.1696

Difference in Two Proportions

-0.0376

Average Proportion

0.1493

Z Test Statistic

-2.1214

Two-Tail Test

Lower Critical Value

-1.9600

Upper Critical Value

1.9600

p-Value

0.0339

Reject the null hypothesis

CI = p1-p2 ± z*se

Standard error of p1-p2= se=

Confidence Interval Estimate

of the Difference Between Two Proportions

Data

Confidence Level

90%

Intermediate Calculations

Z Value

1.645

Std. Error of the Diff. between two Proportions

0.0178

Interval Half Width

0.0293

Confidence Interval

Interval Lower Limit

-0.0669

Interval Upper Limit

-0.0082